I need c,d,e, and f please.

media%2Ff92%2Ff92f315d-6bdf-4601-af16-93
Answer #1

a) The heat of combustion of benzoic acid is -3227 KJ/mol

The equation for the combustion is,

C7H6O2(s) + 7.5 O2(g) -> 7CO2(g) + 3H2O(l), dH° = 3227 kJ

Since 7.5 moles of O2 gas is needed, and 7 moles of CO2 is produced, some pressure-volume work is done, to the calorimeter:

P V = dn * R * T, where dn = (7 - 7.5) = - 0.5 mol
dE = dH - dn R T
= - 3227 - (-0.5*    0.008314 kJ/(K mol) * 298)

= - 3226 kJ/mol (a small correction)

The amount of heat produced by 2.0 g benzoic acid is

q = 2.0 /122.13 mol x 3226 (kJ/mol) = 52.83kJ

Thus, the heat capacity is

C = q / dT = 52.83 / (23.34 -17.84)

C= 52.83 / 5.5 = 9.6 kJ / K.(answer)